1.ONE NUMBER IS 19 LESS THAN THE OTHER AND THEIR SUM IS 65.
2.FIND TWO NUMBERS WHOSE SUM IS 39 AND DIFFERENCE IS 7.
3.IF TWICE NUMBER IS INCREASED BY IN THE RESULT IS 35.

Answer by texttutoring(324)

1. Let x = the number, the other number would be x-19
x + x-19 = 65
2x-19=65
2x=84
x=84/2
x=42
The other number is 42-19=23. So the two numbers are 42 and 23.
2. Similar to the last one...

x+x-7 = 39
2x-7=39
2x=46
x=46/2
x=23
The other number is 16.

I don't understand the 3rd question, sorry.

Hope that helps for the first two!